### v1

# LRT-Diffusion: Calibrated Risk-Aware Guidance for Diffusion Policies

Abstract
[…]
Keywords Offline Reinforcement learning; Diffusion Policies; Likelihood‑ratio Test; Inference; Risk‑aware Guidance; Calibration; Distribution Shift; Q‑guidance; Out‑of‑distribution Detection; D4RL.
1 Introduction
Offline reinforcement learning (RL) aims to learn high‑performing policies from fixed datasets without further environment interaction. A central difficulty is distributional shift: actions proposed by the learned policy can drift away from the behavioral support where value estimates are reliable. Diffusion policies have recently emerged as strong generative decision‑makers for offline RL Janner2022Diffuser : by learning a conditional diffusion model over actions given state, they produce smooth, high‑fidelity samples that respect support better than direct regression. However, how these policies are guided at sampling time remains largely heuristic. Common practices—injecting QQ‑gradients with hand‑tuned schedules and ad‑hoc clipping—lack a statistical notion of risk and offer limited control of the return–shift trade‑off. Unlike prior works, our approach keeps training strictly vanilla (no critic-guided losses) and moves all risk control to inference via a calibrated likelihood-ratio gate (LRT), yielding an interpretable, reproducible risk knob.
Intuition.
Rather than always pulling samples toward the conditional head, we ask at every denoising step: is there enough evidence to move toward the “good‑action” direction for this state? Concretely, we split offline actions into a good subset and a background subset, train a two‑head diffusion model (an unconditional head on all data and a conditional head on good data) with class‑balancing and optional advantage‑based soft weights, and then gate the conditional pull at inference by a calibrated likelihood ratio. The gate is motivated by the Neyman–Pearson test NeymanPearson1933 : a single user knob Type‑I rate $\alpha$ controls the tolerated false activations under $H_{0}$ (“background is correct”); while the hard LRT is UMP at level $\alpha$ under equal covariances, we use a smooth gate in practice for numerical stability and keep $\alpha$ interpretable via calibration on held‑out states.
Method overview.
We introduce LRT‑Diffusion, a risk‑aware, inference‑only sampling scheme for diffusion policies. At each denoising step we make a binary decision between a background prediction and a data‑conditioned prediction. We accumulate evidence during sampling and open the gate only when the evidence is strong, interpolating between the two predictions with a data‑dependent weight. A single threshold is calibrated once so that the empirical false‑activation rate does not exceed a user‑chosen $\alpha$, turning guidance from a fixed push into an evidence‑driven adjustment: weak evidence $\Rightarrow$ stay near the background prior; strong evidence $\Rightarrow$ move decisively toward the conditional policy. Training is unchanged.
Composition with value guidance.
We optionally combine LRT gating with a small critic step: at each denoising update, take a capped step that increases the learned critic at the current proposal (standard decreasing schedule). The gradient can be evaluated near the background proposal, the LRT-gated proposal, or an interpolation; the threshold is calibrated with the same choice, so the Type-I semantics are preserved.
Theoretical guarantees at a glance.
Under equal‑covariance heads, the hard gate(thresholding under LRT assumptions) is the uniformly most powerful test at level $\alpha$; our soft gate (a numerically stable surrogate) retains the threshold’s semantics and improves stability. Beyond testing optimality, we analyze offline‑RL error propagation and show that reducing the policy’s state‑conditional OOD rate—which LRT lowers by construction—tightens a lower bound on true return relative to pure QQ‑guidance when the critic is unreliable off‑support.
Practical picture.
The gate exposes monotone, interpretable knobs: $\alpha$ (risk), $\beta_{\max}$ (max pull), and $\delta$ (sharpness). LRT can be combined with a small critic step evaluated at the background mean, the LRT-gated mean, or a blend; the threshold is calibrated with that same choice so the Type-I semantics are preserved. Empirically on D4RL MuJoCo, LRT tracks the target $\alpha$ and improves the return–OOD Pareto relative to strong $Q$-guided baselines in our implementation.
1.1 Contributions
Calibrated LRT guidance for diffusion policies.
Unlike prior research integrating value or energy guidance into the diffusion model, we cast each denoising step as a likelihood-ratio test between a background head and a “good” head, and calibrate a single threshold $\tau$ on held-out states to bound the trajectory-level Type-I rate at a user-chosen $\alpha$. Training remains vanilla (two heads with class balancing and optional advantage weights).
Labeling-and-weighting recipe for the conditional head.
Rather than fitting a single conditional diffusion model to all (state, action) pairs, as is common in return-dominated diffusion or $Q$-guided training, we split offline actions into a good subset and a background subset using IQL advantages (top-$p$ quantile), and train a two-head diffusion model with class-balancing and advantage-aware soft weights. This improves the quality of the conditional head without extra supervision.  [^1] E.g., Diffuser/Decision‑Diffuser use return conditioning and sampling-time guidance; Diffusion‑QL injects value signals into the training loss, all without a binary good/background head.
Actionable theory with finite-sample and stability guarantees.
Under equal covariances the hard LRT is UMP at level $\alpha$ (Prop. 1); a one-shot Monte-Carlo calibration yields $\mathbb{P}_{H_{0}}(\ell_{\mathrm{cum}}\geq\hat{\tau})\leq\alpha+\varepsilon_{n}$ (Thm. 2). We bound the per-step deterministic drift and show the cumulative LLR is sub-Gaussian with an explicit variance proxy (Sec. 5.2), explaining why calibration is numerically stable.
Return bounds that connect risk control to distribution shift.
With a standard offline error split, we prove a comparison bound for true returns between LRT-guided and $Q$-guided sampling (Prop. 3). A mild monotonicity assumption turns the calibrated level $\alpha$ into a conservative OOD upper bound (Prop. 4), clarifying when and why LRT is preferable to pure $Q$-guidance.
Compatibility with value gradients via matched calibration.
LRT controls risk while the QQ‑step pursues return, and these roles are orthogonal. We show that composing LRT with a small $Q$-step preserves level-$\alpha$ semantics so long as calibration uses the exact deployed sampler (Prop. 2). Thus, risk control (through $\alpha$ and its calibrated $\tau$) is decoupled from return‑seeking (through $\nabla_{a}\hat{Q}$), making the QQ‑update a plug‑and‑play module. The knobs $(\alpha,\beta_{\max},\delta)$ remain monotone and interpretable.
Empirical validation.
On D4RL MuJoCo, LRT-Diffusion honors the target Type-I rate and improves the return–OOD Pareto over strong $Q$-guided baselines in our implementation, with ablations over $\alpha$, $\beta_{\max}$, $\delta$, and gradient anchoring.
2 Related Work
Distribution shift in offline RL and conservative learning.
A central obstacle in offline RL is that policy actions may fall outside the behavior support, where critic estimates are brittle. Conservative algorithms explicitly counter this by penalizing or constraining OOD actions, e.g., BCQ Fujimoto2019BCQ , BEAR Kumar2019BEAR , CQL Kumar2020CQL , and TD3+BC Fujimoto2021TD3BC . IQL Kostrikov2022IQL learns values via expectile regression and performs advantage-weighted improvement without explicit behavior cloning. Our work is complementary: we keep training unchanged (e.g., use IQL advantages only to label/weight “good” actions) and control OOD at sampling time via a calibrated gate.
Diffusion policies for decision-making and their guidance.
Diffusion models have been adopted for control and planning (e.g., Diffuser Janner2022Diffuser ). To improve returns, many systems apply Q-guidance—adding an action-space step along $\nabla_{a}\hat{Q}(s,a)$ with hand-tuned schedules and clips—conceptually analogous to guidance in image diffusion (e.g., classifier-based guidance Dhariwal2021GuidedDiffusion and classifier-free guidance Ho2022CFG ). However, such heuristics lack a statistical notion of risk: there is no global control on the probability of “falsely” pulling away from the background prior.
Risk-aware control and calibration.
Risk-sensitive and conservative offline RL methods (e.g., behavior-regularized or pessimistic objectives) encode risk in the training loss by penalizing value estimates or constraining policy deviation from the dataset (Fujimoto2019BCQ, ; Wu2019BRAC, ; Kumar2020CQL, ; Nair2020AWAC, ; Peng2019AWR, ; Kostrikov2022IQL, ). A complementary line uses distribution-free calibration to turn data-driven thresholds into finite-sample guarantees (e.g., conformal prediction) (Vovk2005ALRW, ; Angelopoulos2023ConformalSurvey, ). In diffusion-based decision making, sampling is typically guided heuristically (e.g., $Q$-gradient pushes with hand-tuned schedules/clipping) without an explicit statistical notion of risk. Our approach bridges these threads: we cast each reverse step as a simple-vs-simple test between an unconditional (background) head and a conditional (good) head, accumulate a log-likelihood ratio (LLR), and calibrate one threshold on held-out states so that the empirical Type-I rate under $H_{0}$ does not exceed a user-chosen $\alpha$. Under equal covariances—satisfied by our two-head design—the hard likelihood-ratio test is uniformly most powerful (Neyman–Pearson) (NeymanPearson1933, ), and we use a smooth gate for numerical stability while preserving the same $\alpha$-semantics via calibration. This gives a statistically grounded alternative to heuristic mixing: guidance becomes evidence-driven with an explicit level-$\alpha$ risk budget at inference time, without changing training.
Hypothesis testing and likelihood ratios in RL
Statistical testing has been used in RL primarily for off-policy evaluation and safe policy improvement, not for stepwise guidance inside a generative sampler. High-confidence OPE Thomas2015HCOPE derives concentration-based bounds to accept/reject policy deployment with user-specified risk, while SPIBB Laroche2019SPIBB ensures safe improvement over a baseline via count-based constraints rather than explicit tests. In safety-constrained RL, trust-region and constraint-violation controls (e.g., CPO Achiam2017CPO ) target reliable deployment but do not implement likelihood-ratio gating during action generation. Classical sequential tests such as Wald’s SPRT Wald1945SPRT have inspired sequential decision criteria, yet we are not aware of prior work that uses a calibrated likelihood-ratio test as an inference-time gate to steer diffusion-policy denoising steps. Our method fills this gap by turning denoising guidance into an evidence-controlled, level-$\alpha$ process.
Advantage-based labeling and weighting.
For higher return actions, advantage-weighted and critic-regularized schemes (RWR Peters2007RWR , AWR Peng2019AWR , AWAC Nair2020AWAC ) are proposed. We use IQL advantages to (i) define a top-$p$ “good” subset for the conditional head and (ii) optionally apply a temperature-controlled soft weight on positives, while keeping a background head trained on all data. This two-head setup strengthens the conditional signal without altering the base training pipeline, and remains compatible with our calibrated LRT at inference. Our two-head design enforces equal covariances by construction, which simplifies the LLR and aligns with the NP test.
3 Background
Offline RL is given a fixed dataset $\mathcal{D}=\{(s^{(i)},a^{(i)},r^{(i)},s^{(i)^{\prime}})\}_{i=1}^{N}$, collected by an unknown behavior policy $\pi$, and aims to learn a high-return policy while controlling out-of-distribution risk. In the dataset, $s\in\mathbb{R}^{d_{s}}$ is the current state, $a\in\mathbb{R}^{d_{a}}$ is the action, $r\in\mathbb{R}$ is the reward, $s^{\prime}$ is the next state. We consider the process as a discounted Markov Decision Process with continuous state/action spaces.
3.1 IQL Advantages and “Good vs. Background” Labels
For each pair of state and action $(s,a)\in\mathcal{D}$, we train an IQL-style critic Kostrikov2022IQL $(\hat{Q}(s,a),\hat{V}(s))$ on standardized inputs and define an advantage $A(s,a)$ as:
$A(s,a)\;=\;\hat{Q}(s,a)-\hat{V}(s),$
which quantifies the value contributed by action $a$ at state $s$ relative to the state’s baseline.
Let $\kappa$ be the (1p)(1-p)-quantile of $\{A(s_{i},a_{i})\}_{i=1}^{N}$. We label each pair of state-action data by
$c\;=\;\mathbf{1}\{A(s,a)\geq\kappa\}\in\{0,1\}.$
Pairs with c=1c{=}1 are considered good, as their advantages fall within the top (1p)(1-p) fraction of the dataset, while those with c=0c{=}0 form the non-advantage subset. In practice, we set $p=0.2$, allocating approximately $20\%$ of the data as good examples.
3.2 Diffusion Policies with Two Heads
We train a diffusion policy with two heads: an unconditional head trained on all state–action pairs $\mathcal{D}$, and a conditional head trained only on the high-advantage subset (c=1c{=}1) we labeled before. Formally, let $\mathcal{D}_{\mathrm{all}}:=\mathcal{D}$ and $\mathcal{D}_{\mathrm{good}}:=\{(s,a,r,s^{\prime})\in\mathcal{D}:c=1\}$. The unconditional head learns broad dataset coverage from $\mathcal{D}_{\mathrm{all}}$, while the conditional head specializes to high-advantage behavior using $\mathcal{D}_{\mathrm{good}}$.
3.2.1 Two heads and notation
We use a shared backbone $\phi_{\theta}(s,a_{t},t)$ with two output branches (“heads”): an unconditional/background head trained on $\mathcal{D}_{\mathrm{all}}$, and a conditional/good head trained on $\mathcal{D}_{\mathrm{good}}$. Each head maps features to a DDPM-style noise prediction,
$\hat{\varepsilon}_{u}\;=\;h_{u,\theta}\!\big(\phi_{\theta}(s,a_{t},t)\big),\qquad\hat{\varepsilon}_{c}\;=\;h_{c,\theta}\!\big(\phi_{\theta}(s,a_{t},t)\big),$ (1)
where the subscripts $u$ and $c$ will consistently denote unconditional and conditional quantities, respectively.
3.2.2 DDPM parameterization
We adopt the predict-$\varepsilon$ parameterization Ho2020DDPM ; Nichol2021ImprovedDDPM with schedule $\{\alpha_{t},\bar{\alpha}_{t}\}_{t=1}^{T}$, where $\bar{\alpha}_{t}=\prod_{\tau=1}^{t}\alpha_{\tau}$. Given the current latent action $a_{t}$ at step $t$, a noise prediction $\hat{\varepsilon}$ induces a Gaussian reverse kernel with shared covariance $\sigma_{t}^{2}I$ and mean
[…]
We define the head-specific means
$\mu_{u}(t,s,a_{t}):=\mu(\hat{\varepsilon}_{u};t,s,a_{t}),\qquad\mu_{c}(t,s,a_{t}):=\mu(\hat{\varepsilon}_{c};t,s,a_{t}),$
and use the same $\sigma_{t}>0$ from the diffusion schedule for both heads. This yields two reverse kernels for the next-step action:
$a_{t-1}\mid a_{t},s~\sim~\mathcal{N}\!\big(\mu_{u}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),\quad a_{t-1}\mid a_{t},s~\sim~\mathcal{N}\!\big(\mu_{c}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),$
where $\mathcal{N}(\mu,\Sigma)$ is Gaussian with mean $\mu$ and variance $\Sigma$.
After training, the parameters $\theta$ are frozen, so $\mu_{u}(t,s,a_{t})$ and $\mu_{c}(t,s,a_{t})$ are deterministic functions of $(t,s,a_{t})$ (with the same $\sigma_{t}$). We will interpret a single reverse step through these two proposals in the next subsection.
Notation clarification.
We write dataset indices as parenthesized superscripts, e.g., $a^{(i)}$ is the ithi-th action in the dataset $\mathcal{D}$. We reserve subscripts $t$ for diffusion steps, e.g., $a_{t}$ is the latent at step $t$ and $a_{0}$ is the final action..
3.3 Hypotheses for a Reverse Step
With the trained parameters frozen, each reverse step $t$ and conditioning state–latent pair $(s,a_{t})$ yields two Gaussian proposals for the next latent action $a_{t-1}$:
$H_{0}~(\text{background}):~~a_{t-1}\mid a_{t},s~\sim~\mathcal{N}\!\big(\mu_{u}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),$
$H_{1}~(\text{good}):~~a_{t-1}\mid a_{t},s~\sim~\mathcal{N}\!\big(\mu_{c}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),$
where $\mu_{u}$ and $\mu_{c}$ are the head-specific means defined in the previous subsection 3.2. Intuitively, $H_{0}$ favors broad, background behavior supported by the entire dataset, while $H_{1}$ emphasizes high-advantage behavior learned from the c=1c{=}1 subset.
For compactness, denote the head-induced one-step densities by
$p_{c}(a_{t-1}\mid a_{t},s,t):=$ $\mathcal{N}\!\big(a_{t-1};\,\mu_{c}(t,s,a_{t}),\,\sigma_{t}^{2}I\big),$ (2)
$p_{u}(a_{t-1}\mid a_{t},s,t):=$ $\mathcal{N}\!\big(a_{t-1};\,\mu_{u}(t,s,a_{t}),\,\sigma_{t}^{2}I\big).$ (3)
3.4 Likelihood-Ratio View of a Reverse Step
Unknown labels $\Rightarrow$ a trajectory-level test.
At inference the label $c$ is unknown. Given two simple trajectory models $p_{u}(a_{T:1}\mid s)$ (background) and $p_{c}(a_{T:1}\mid s)$ (good), the Neyman–Pearson (NP) lemma implies that the level-$\alpha$ test that rejects $H_{0}$ when the cumulative log–likelihood ratio (LLR) exceeds a threshold is uniformly most powerful (UMP). We therefore base our gate on the trajectory statistic
$\ell_{\mathrm{cum}}(a_{T:1})\;:=\;\log\frac{p_{c}(a_{T:1}\mid s)}{p_{u}(a_{T:1}\mid s)}.$
How to compute $\ell_{\mathrm{cum}}$: step-wise LLRs.
With the reverse chain factorization in DDPM, for $i\in\{u,c\}$
$p_{i}(a_{T:1}\mid s)=q(a_{T})\prod_{t=1}^{T}p_{i}(a_{t-1}\mid a_{t},s,t),$
so the trajectory LLR decomposes into a sum of step-wise terms:
$\ell_{\mathrm{cum}}(a_{T:1})~=~\sum_{t=1}^{T}\ell_{t},\qquad\ell_{t}~:=~\log\frac{p_{c}(a_{t-1}\mid a_{t},s,t)}{p_{u}(a_{t-1}\mid a_{t},s,t)}.$ (4)
Head-induced one-step densities and quadratic form.
From Sec. 3.3, each head induces a Gaussian kernel with shared covariance. Hence the step-wise LLR admits the shared-variance simplification
$\ell_{t}~=~\frac{1}{2\sigma_{t}^{2}}\Big(\|a_{t-1}-\mu_{u}(t,s,a_{t})\|_{2}^{2}-\|a_{t-1}-\mu_{c}(t,s,a_{t})\|_{2}^{2}\Big),$ (5)
i.e., a linear discriminant in $a_{t-1}$ and cheap to evaluate.
Decision rule for assigning labels.
The NP test uses $\ell_{\mathrm{cum}}$ with a calibrated threshold $\tau$ to control the level-$\alpha$ Type-I rate. We implement a smooth gate that monotonically approximates the NP decision and calibrate $\tau$ under $H_{0}$; details are given in Sec. 4.
4 Method: LRT-Diffusion
We split offline actions by an advantage threshold into a good subset and a background subset. We then train a vanilla two‑head diffusion policy (DDPM $\epsilon$‑prediction Ho2020DDPM ): an unconditional head on all data and a conditional head on the good subset, with class‑balancing and optional advantage‑based soft weights; no value/energy‑guided losses are added. At inference, each reverse step is treated as a binary test between heads: we accumulate a cumulative log‑likelihood ratio (LLR; Sec. 3.4) and apply a calibrated logistic gate to interpolate the mean,
$\mu_{\text{LRT},t}\;=\;\mu_{u}(t,s,a_{t})\;+\;\beta_{t}\,\big(\mu_{c}(t,s,a_{t})-\mu_{u}(t,s,a_{t})\big),$ (6)
then sample $a_{t-1}\sim\mathcal{N}(\mu_{\text{LRT},t},\sigma_{t}^{2}I)$. The threshold $\tau$ is calibrated under $H_{0}$ (background) to meet a user‑chosen Type‑I rate $\alpha$.Below is the end-to-end pipeline:
  1. 1.
[…]
  2. 2.
    Compute advantages and label top-$p$ pairs, see §4.1;
  3. 3.
    Train a two-head diffusion model on $\epsilon$-prediction, see §4.2;
  4. 4.
    Calibrate a single threshold $\tau$ under $H_{0}$, see §4.4;
  5. 5.
    At inference, accumulate a cumulative LLR and gate the conditional pull, see §4.5;
4.1 Good‑vs‑background Labeling
We rank state–action pairs by the IQL advantage $A(s,a)=\hat{Q}(s,a)-\hat{V}(s)$ computed on standardized inputs, and mark the top‑p global quantile as “good.” Although one could define state‑wise thresholds, the advantage already subtracts a state‑dependent baseline $\hat{V}(s)$, making A comparable across states in practice. Moreover, global top‑p is a non‑parametric surrogate of advantage‑weighted learning (e.g., AWR/AWAC), corresponding to a small‑temperature limit without tuning an extra temperature. Practically, offline datasets have highly uneven coverage—many states appear once or with only a handful of actions—so per‑state ranking/thresholding is statistically brittle (often ill‑posed); a single global threshold is more stable and reproducible. Crucially, our risk control via LRT is orthogonal to this choice: once the two‑head model is trained, Type‑I error is calibrated at inference regardless of how the conditional head’s subset was selected.
4.2 Two-head Training
Under the two-head diffusion policy in Sec. 3.2, we optimize the $\epsilon$-prediction loss (refer to Eq. 1) with per-sample weights
$\mathcal{L}=\tfrac{1}{B}\sum_{i=1}^{B}\tilde{w}_{i}\,\big\|\hat{\epsilon}_{\phi}(a_{t,i},t,\tilde{s}_{i})-\epsilon_{i}\big\|_{2}^{2},\quad\tilde{w}_{i}=\frac{w^{\text{cb}}(c_{i};\hat{\rho})\cdot u_{i}}{\frac{1}{B}\sum_{j=1}^{B}w^{\text{cb}}(c_{j};\hat{\rho})\cdot u_{j}},$
[…]
$w^{\mathrm{cb}}(c;\hat{\rho})=\frac{\mathbf{1}\{c=1\}}{2\hat{\rho}+\varepsilon}+\frac{\mathbf{1}\{c=0\}}{2(1-\hat{\rho})+\varepsilon},$ (7)
$u_{i}=1+\mathbf{1}\{c_{i}=1\}\,\min\!\Big\{\max\!\big(0,\tfrac{A_{i}-\kappa}{\tau_{A}}\big),\,u_{\max}-1\Big\}.$ (8)
Multiplying $w^{\text{cb}}$ and $u_{i}$ separates roles: the former fixes class imbalance between conditional and unconditional datasets, while the latter reallocates mass within c=1c{=}1 without changing the effective class ratio; the batch normalization of $\tilde{w}_{i}$ stabilizes the step size.
4.3 Evidence-gated Sampler
4.3.1 Motivation
1D evidence direction.
From Eq. 5, under equal covariances the one-step LLR has constant gradient $\nabla_{a_{t-1}}\ell_{t}=\Sigma_{t}^{-1}(\mu_{c}-\mu_{u})$; among unit directions $u$ the directional derivative $\langle u,\nabla\ell_{t}\rangle$ is maximized by $u\parallel\Sigma_{t}^{-1}(\mu_{c}-\mu_{u})$ (isotropic: $u\parallel\mu_{c}-\mu_{u}$). We therefore restrict the reverse-step mean to the 1D ray $\mu_{u}+\beta(\mu_{c}-\mu_{u}),$ and let the scalar gate $\beta$ depend monotonically on the cumulative evidence $\ell_{\mathrm{cum}}$.
Why not always use $\mu_{c}$?
Always setting $a_{t-1}{=}\mu_{c}{+}\sigma_{t}z_{t}$ is equivalent to always accepting $H_{1}$ (i.e., $\alpha{=}1$), which (i) removes any control on false activations of the conditional head and thus increases off-support mass in offline RL, and (ii) couples performance to critic/label errors where the model is most brittle. Our gate opens only when there is sufficient evidence (large $\ell_{\mathrm{cum}}$), yielding a calibrated budget on false activations (Prop. 2 and Thm. 2). Section 5.3 further shows that, when off-support errors dominate ($\varepsilon_{\rm out}\!\gg\!\varepsilon_{\rm in}$), reducing the state-conditional OOD rate via a smaller $\alpha$ tightens a lower bound on the true return (Prop. 3).
4.3.2 Gate Selection
Hard gate (UMP at level-$\alpha$).
Let $\tau$ be a threshold. The hard likelihood-ratio test accepts $H_{1}$ when $\ell_{\mathrm{cum}}\geq\tau$ and rejects otherwise. Equivalently, with a cap $\beta_{\max}\!\in[0,1]$,
$\beta_{t}^{\text{h}}\;=\;\beta_{\max}\,\mathbf{1}\{\ell_{\mathrm{cum}}\geq\tau\},\quad a_{t-1}\;=\;\mu_{u}\,+\,\beta_{t}^{\text{h}}(\mu_{c}-\mu_{u})\,+\,\sigma_{t}z_{t}.$ (9)
Under equal covariances and the joint factorization of the reverse chain, this hard test is uniformly most powerful (UMP) among all level-$\alpha$ tests (Prop. 1).
Soft gate (stable surrogate) and its hard-limit.
For numerical stability we use a logistic surrogate
$\beta_{t}\;=\;\beta_{\max}\,\sigma\!\Big(\tfrac{\ell_{\mathrm{cum}}-\tau}{\delta}\Big),\quad a_{t-1}\;=\;\mu_{u}\,+\,\beta_{t}(\mu_{c}-\mu_{u})\,+\,\sigma_{t}z_{t},$ (10)
where $z_{t}\!\sim\!\mathcal{N}(0,I)$ and $\delta{>}0$ controls sharpness (smaller $\delta$ $\to$ sharper switch). As $\delta\!\downarrow\!0$, the soft rule converges to the hard gate both pointwise and in trajectory law (Lemma 1). Thus, the $\alpha,\tau,\beta_{t}$ can be seen as monotone, interpretable risk knobs. The resulting sampler is summarized in Alg. 2.
4.4 Calibration
Fix $(\beta_{\max},\delta)$ and freeze the sampler in Alg.2. On a held-out state set matched to deployment, simulate reverse chains under $H_{0}$ using the same sampler and collect the realized $\ell_{\mathrm{cum}}$; set
$\hat{\tau}=\mathrm{Quantile}_{1-\alpha}\big\{\ell_{\mathrm{cum}}^{(i)}\big\}.$
Calibrating with the exact deployment sampler preserves the level-$\alpha$ semantics (Prop. 2). By the Dvoretzky–Kiefer–Wolfowitz bound Dvoretzky1956DKW ; Massart1990 , with probability $\geq 1-\eta$ over $n$ calibration draws,
$\mathbb{P}_{H_{0}}^{\text{(sampler)}}\!\big(\ell_{\mathrm{cum}}\geq\hat{\tau}\big)\;\leq\;\alpha+\sqrt{\tfrac{1}{2n}\log\tfrac{2}{\eta}},$
giving a finite-sample guarantee on the realized Type-I rate (Thm. 2). See Alg. 1 for the Monte Carlo procedure.
4.5 Inference
We calibrate $\tau$ once per (task, model, gate hyperparameters) on a held-out state set, using the same frozen sampler as in Alg. 2 (and Alg. 1 for the Monte-Carlo procedure). Thereafter, all rollouts simply call Alg. 2 with this fixed $\hat{\tau}$; no re-calibration is needed unless $(\beta_{\max},\delta)$ or the optional $Q$-composition change. The returned $a_{0}$ is on the standardized scale and is mapped to the environment as $a_{\mathrm{env}}=a_{0}\odot\mathrm{std}_{a}+\mathrm{mean}_{a}$ (then clipped to action bounds).
Algorithm 1 Calibration under $H_{0}$ to obtain $\tau$ (pre-inference sampling)
1:Frozen policy $\theta$, steps $T$, risk $\alpha$, gate $(\beta_{\max},\delta)$, noises $\{\sigma_{t}\}$, states $\mathcal{S}_{\rm cal}$, budget $n$
2:Threshold $\tau$
[…]
4:for $k=1..K_{\max}$ do
5:  $\mathcal{L}\leftarrow\emptyset$
6:  while $|\mathcal{L}|<n$ do
7:   Sample $s,a_{T}\sim\mathcal{N}(0,I)$; $\ell\leftarrow 0$
8:   for $t=T..1$ do
9:     $(\mu_{u},\mu_{c})\leftarrow\mu_{\theta}(a_{t},s)$
10:     $\beta_{t}\leftarrow\beta_{\max}\sigma((\ell-\tau)/\delta)$(soft-gate)
11:     $a_{t-1}\sim\mathcal{N}(\mu_{u}+\beta_{t}(\mu_{c}-\mu_{u}),\sigma_{t}^{2}I)$
12:     $\ell\leftarrow\ell-\frac{\|a_{t-1}-\mu_{c}\|^{2}-\|a_{t-1}-\mu_{u}\|^{2}}{2\sigma_{t}^{2}}$
13:   end for
14:   $\mathcal{L}\leftarrow\mathcal{L}\cup\{\ell\}$
15:  end while
16:  $\tau\leftarrow{\rm Quantile}_{1-\alpha}(\mathcal{L})$
17:end for
18:return $\tau$
Algorithm 2 LRT-Guided Inference
1:State $s$, steps $T$, $\tau$, $(\beta_{\max},\delta)$, noises $\{\sigma_{t}\}$
2:Action $a_{0}$
[…]
4:for $t=T..1$ do
5:  $(\mu_{u},\mu_{c})\leftarrow\mu_{\theta}(a_{t},s)$
6:  $\beta_{t}\leftarrow\beta_{\max}\sigma((\ell-\tau)/\delta)$
7:  $a_{t-1}\sim\mathcal{N}(\mu_{u}+\beta_{t}(\mu_{c}-\mu_{u}),\sigma_{t}^{2}I)$
8:  $\ell\leftarrow\ell-\frac{\|a_{t-1}-\mu_{c}\|^{2}-\|a_{t-1}-\mu_{u}\|^{2}}{2\sigma_{t}^{2}}$
9:  $a_{t}\leftarrow a_{t-1}$ (optionally use Eq. 11)
10:end for
11:return $a_{0}$
4.6 Composition with Value Gradients
To cleanly separate risk control (via the LRT gate) from return seeking, we optionally add a small critic step that mirrors common practice in diffusion-policy samplers. This step uses the action-space gradient of a learned critic (not the diffusion score) with a simple schedule and clipping:
$a_{t-1}\ \leftarrow\ a_{t-1}\ +\ \lambda_{t}\,\sigma_{t}^{2}\,\nabla_{a}\hat{Q}_{\theta}(s,a)\big|_{a=a_{c}},$ (11)
where $a_{c}\in\{\mu_{u},\ \mu_{\text{LRT},t},\ (1-\rho)\mu_{u}+\rho\,\mu_{\text{LRT},t}\},$ and $\mu_{\text{LRT},t}=\mu_{u}+\beta_{t}(\mu_{c}-\mu_{u})$ is the LRT-gated mean. We use a light, hand-tuned schedule (e.g., $\lambda_{t}\propto\sigma_{t}$) and gradient clipping to keep updates stable.
We use the evaluation center $a_{c}=\mu_{\text{LRT},t}$ as evidence opens the gate, suggesting the gradient is closer to the good direction. For more evaluation center choice discussion, refer Appendix D.1
5 Theory and Properties
5.1 Gate Selection: UMP vs. Stability
In this part, we discuss the properties of hard and soft gates. For the full mathematical proof, see Appendix A.
Proposition 1 (Neyman–Pearson optimality).
For the reverse chain conditioned on $s$ with two simple hypotheses follows the two-head reverse model of Sec. 3.4. Then the Neyman-Pearson test that rejects $H_{0}$ when $\ell_{\mathrm{cum}}\geq\tau$ is uniformly most powerful among all level-$\alpha$ tests.
[…]
Assume the two-head reverse model of Sec. 3.4 and a soft gate denoted in Sec. 4.3.2. At step $t$, the proposal mean is $\mu_{t}^{\mathrm{soft}}=\mu_{u}+\beta_{t}(\mu_{c}-\mu_{u})$ and the reverse variance is $\sigma_{t}^{2}I$. Fix $(\tau,\beta_{\max})$. As $\delta\to 0$, we have pointwise
$\mu_{t}^{\mathrm{soft}}\;\to\;\mu_{t}^{\mathrm{hard}}:=\mu_{u}+\beta_{\max}\,\mathbf{1}\{\ell_{\mathrm{cum}}\geq\tau\}\,(\mu_{c}-\mu_{u}),$
and the trajectory law induced by the soft-gated sampler converges weakly to that of the hard-gated sampler.
Proposition 2 (Calibrated semantics under the deployment sampler).
Fix sampler hyperparameters $(\beta_{\max},\delta)$ (and any deterministic $Q$-composition). Let $\hat{\tau}$ be the empirical $(1-\alpha)$ quantile of $\ell_{\mathrm{cum}}$ computed from i.i.d. rollouts under $H_{0}$ with the same frozen sampler. Then the realized false-activation rate at deployment satisfies
$\mathbb{P}_{H_{0}}^{(\text{sampler})}\!\big(\ell_{\mathrm{cum}}\geq\hat{\tau}\big)\approx\alpha,$
up to finite-sample fluctuations (see Thm. 2 for a DKW bound).
Discussion.
The hard gate is UMP at level $\alpha$ (Prop. 1); the soft gate (Lemma 1) provides a numerically stable approximation with the same level-$\alpha$ semantics when $\tau$ is calibrated under the exact deployment sampler (Prop. 2).
5.2 Finite-sample Calibration and Stability
Let $F_{0}$ be the CDF of the cumulative LLR $\ell_{\mathrm{cum}}$ under $H_{0}$ for the frozen sampler (soft gate and, if present, the fixed $Q$-step). Given i.i.d. calibration draws $\ell^{(1)},\dots,\ell^{(n)}$ and the empirical CDF $\widehat{F}_{n}$, define the plug-in quantile $\hat{\tau}=\inf\{x:\widehat{F}_{n}(x)\geq 1-\alpha\}$.
5.2.1 Finite-sample Guarantee
Theorem 2 (Calibration accuracy via DKW Dvoretzky1956DKW ; Massart1990 ).
For any $\zeta\in(0,1)$, with probability at least $1-\zeta$ over the calibration sample,
$\mathbb{P}_{H_{0}}^{\text{(sampler)}}\!\big(\ell_{\mathrm{cum}}\geq\hat{\tau}\big)~\leq~\alpha+\varepsilon_{n},\qquad\varepsilon_{n}=\sqrt{\tfrac{1}{2n}\log\tfrac{2}{\zeta}}.$
Calibration budget. Rearranging the bound gives a one-line rule: to guarantee $\mathbb{P}_{H_{0}}^{(\text{sampler})}(\ell_{\mathrm{cum}}\!\geq\!\hat{\tau})\leq\alpha+\varepsilon$ with confidence at least $1-\zeta$, it suffices to set sample size $n\;\geq\;\frac{1}{2\varepsilon^{2}}\,\log\!\frac{2}{\zeta}.$ See Appendix B.1 for proof.
5.2.2 Stability
Lemma 3 (Deterministic displacement bound).
At reverse step $t$, write $\Delta\mu_{t}:=\mu_{c,t}-\mu_{u,t}$ and let $g_{t}$ be the clipped critic gradient with $\|g_{t}\|\leq G$. Our update has deterministic mean
$m_{t}\;=\;\mu_{u,t}\;+\;\beta_{t}\,\Delta\mu_{t}\;+\;\lambda_{t}\,\sigma_{t}^{2}\,g_{t},\qquad 0\leq\beta_{t}\leq\beta_{\max},\ \ 0\leq\lambda_{t}\leq\lambda_{\max}.$
Hence the per-step deterministic displacement from the background anchor is bounded by
$\|m_{t}-\mu_{u,t}\|~\leq~\beta_{\max}\,\|\Delta\mu_{t}\|\;+\;\lambda_{\max}\,\sigma_{t}^{2}\,G\;=:\;B_{t}.$ (12)
If, additionally, $\|\Delta\mu_{t}\|\leq D$ (e.g. clamp on $\Delta\mu_{t}$) and $\sigma_{t}^{2}\leq S^{2}$, then
$\|m_{t}-\mu_{u,t}\|\leq\beta_{\max}D+\lambda_{\max}S^{2}G\;=:\;B_{\text{step}}$
for all $t$, and the cumulative deterministic deviation from the background chain across $T$ steps is at most $\sum_{t=1}^{T}B_{t}\leq T\,B_{\text{step}}$.
Thus, the deterministic component of the sampler admits a uniform, hyperparameter‑controlled drift bound both per step and over the full reverse trajectory; see Appendix B.2 for the detailed proof.
Implication for one-step LLR fluctuations.
Under equal covariances, the one-step LLR increment admits the identity
$\Delta\ell_{t}~=$~ $\underbrace{\frac{\Delta\mu_{t}^{\top}\big(m_{t}-\tfrac{\mu_{c,t}+\mu_{u,t}}{2}\big)}{\sigma_{t}^{2}}}_{\text{deterministic part}}\;+\;\underbrace{\frac{\Delta\mu_{t}^{\top}z_{t}}{\sigma_{t}}}_{\text{zero-mean Gaussian}},$
where $z_{t}\sim\mathcal{N}(0,I)$. Therefore, conditionally on $(m_{t},\mu_{u,t},\mu_{c,t})$,
$\big|\mathbb{E}[\Delta\ell_{t}]\big|$ $~\leq~\frac{\|\Delta\mu_{t}\|}{\sigma_{t}^{2}}\,\Big(B_{t}+\tfrac{1}{2}\|\Delta\mu_{t}\|\Big),$
$\mathrm{Var}(\Delta\ell_{t})$ $=~\frac{\|\Delta\mu_{t}\|^{2}}{\sigma_{t}^{2}}.$
In particular, if $\|\Delta\mu_{t}\|\leq D$ and $\sigma_{t}\geq\sigma_{\min}>0$, then $\mathrm{Var}(\Delta\ell_{t})\leq D^{2}/\sigma_{\min}^{2}$ and $\big|\mathbb{E}[\Delta\ell_{t}]\big|\leq\frac{D}{\sigma_{\min}^{2}}\big(B_{\text{step}}+\tfrac{D}{2}\big)$, so $\ell_{\mathrm{cum}}=\sum_{t}\Delta\ell_{t}$ is sub-Gaussian with variance proxy $\sum_{t}\|\Delta\mu_{t}\|^{2}/\sigma_{t}^{2}\leq T\,D^{2}/\sigma_{\min}^{2}$. This quantifies that our gate and gradient clipping keep both the magnitude of mean shifts and the variance of the accumulated evidence controlled. See proofs on Appendix B.3 and  B.4.
5.3 Distributional Control and Return Bounds
Let $\mathcal{S}(s)$ denote the dataset action support at state $s$, and define the state-conditional OOD rate of a policy $\pi$ by
$\eta(\pi\,|\,s)=\Pr_{a\sim\pi(\cdot|s)}[a\notin\mathcal{S}(s)],\quad\eta(\pi)=\mathbb{E}_{s\sim d_{\text{eval}}}\![\eta(\pi\,|\,s)].$
Let $\hat{Q}$ be a learned critic and $Q^{\text{true}}$ the environment action-value. [^2] Any fixed evaluation state distribution $d_{\text{eval}}$ may be used; in our experiments it is the dataset state marginal. Assume the standard offline error split
$\varepsilon_{\rm in}:=\sup_{a\in\mathcal{S}(s)}|\hat{Q}(s,a)-Q^{\text{true}}(s,a)|,\quad\varepsilon_{\rm out}:=\sup_{a\notin\mathcal{S}(s)}|\hat{Q}(s,a)-Q^{\text{true}}(s,a)|,$
and that $Q^{\text{true}}$ is $L$-Lipschitz in $a$. Intuitively, $\nu:=\varepsilon_{\rm out}-\varepsilon_{\rm in}\geq 0$ in offline RL due to extrapolation error. Additionally, write $\pi_{\rm LRT}$ for the LRT-gated policy and $\pi_{Q}$ for pure $Q$-guided sampling, with $a_{\rm LRT}\!\sim\!\pi_{\rm LRT}(\cdot|s)$ and $a_{Q}\!\sim\!\pi_{Q}(\cdot|s)$.
Proposition 3 (Return comparison under offline errors).
Under the assumption above, let
$\Delta_{Q^{true}}:$ $=\mathbb{E}_{s}[Q^{true}(s,a_{LRT})-Q^{true}(s,a_{Q})],$
$\Delta_{\hat{Q}}$: $=\mathbb{E}_{s}[\hat{Q}(s,a_{LRT})-\hat{Q}(s,a_{Q})],$
the following inequality holds:
$\boxed{\Delta_{Q^{true}}\geq\Delta_{\hat{Q}}-2\varepsilon_{\rm in}-\nu\,\left(\eta(\pi_{Q})+\eta(\pi_{\rm LRT})\right).}$
Sketch.
Decompose expectations into on-support and OOD parts. On-support deviations are bounded by $\varepsilon_{\rm in}$, off-support by $\varepsilon_{\rm out}$. Refer to Appendix  C.1 for complete proof. ∎
Assumption (Monotone support w.r.t. gating). To relate $\eta(\pi_{\rm LRT})$ to the calibrated level $\alpha$, we use a mild monotonicity assumption. With the background head fixed, opening an LRT gate at any step does not decrease the probability that the final $a_{0}$ is outside $\mathcal{S}(s)$; if all $T$ gates reject, the final $a_{0}$ lies in $\mathcal{S}(s)$ with high probability.
Proposition 4 (From level-$\alpha$ to an OOD bound).
Under the assumption above, if each step’s gate has size at most $\alpha$ under $H_{0}$ and the sampler uses $T$ gated steps, then
$\eta(\pi_{\rm LRT})~\leq~\Pr(A)~\leq~1-(1-\alpha)^{T}~\approx~\alpha T\quad(\alpha T\ll 1).$
A is when at least one gate opens under $H_{0}$.
Remarks. The bound follows from a union/Bonferroni argument (or Ville’s inequality Ville1939 ) for the activation events. It formalizes the role of $\alpha$ as a risk knob: smaller $\alpha$ $\Rightarrow$ fewer/softer activations $\Rightarrow$ lower OOD rate, which—via Prop. 3—tightens a lower bound on the true return relative to pure $Q$-guidance when $\varepsilon_{\rm out}\!\geq\!\varepsilon_{\rm in}$.
5.4 A sufficient condition for LRT to dominate Q
Combining Prop. 3 and Prop. 4 yields
$\Delta_{Q^{\rm true}}\ \geq\ \Delta_{\hat{Q}}\ -\ 2\varepsilon_{\rm in}\ -\ \nu\big(\alpha T+\eta(\pi_{Q})\big).$
Hence, if
$\boxed{\alpha\ \leq\ \alpha_{\max}\ :=\ \frac{\Delta_{\hat{Q}}\ -\ 2\varepsilon_{\rm in}\ -\ \nu\,\eta(\pi_{Q})}{\nu\,T}\,,}$
then $\mathbb{E}_{s}[Q^{\rm true}(s,a_{\rm LRT})]\geq\mathbb{E}_{s}[Q^{\rm true}(s,a_{Q})]$. Moreover, $\alpha_{\max}>0$ if and only if $\Delta_{\hat{Q}}>2\varepsilon_{\rm in}+\nu\,\eta(\pi_{Q})$.
The bound is sufficient (and conservative): it uses the union proxy $\eta(\pi_{\rm LRT})\!\leq\!\alpha T$ and the critic gap $\Delta_{\hat{Q}}$, which may be biased off–support. Nevertheless it yields actionable levers: (i) decreasing $T$ (late–step gating) or $\beta_{\max}$ reduces $\nu$ and enlarges the feasible range; (ii) making $\eta(\pi_{Q})$ small (a conservative anchor for the QQ–step) increases $\alpha_{\max}$; (iii) evidence–tied guidance near $\mu_{\text{LRT}}$ can raise $\Delta_{\hat{Q}}$. In the favorable regime where $\eta(\pi_{Q})$ is small, $\alpha_{\max}$ grows as $\alpha_{\max}=(\Delta_{\hat{Q}}-2\varepsilon_{\rm in}-\nu\,\eta(\pi_{Q}))/(\nu T)$, so moderate $\alpha$ can still be certified. In practice we treat $\widehat{\alpha}_{\max}$ as a diagnostic and sweep $\alpha$ on a log grid; the selected $\alpha$ is the knee of the return–risk curve (Fig. 1), while the certificate provides a sanity upper bound. For further discussion, see Appendix C.5
6 Experiments
We empirically study whether LRT guidance delivers a calibrated, interpretable risk knob at inference and improves the return–OOD trade‑off over standard QQ‑guided sampling. Unless otherwise noted, diffusion training is vanilla (Sec. 4.2); all risk control is applied only at inference via the calibrated LRT gate.
6.1 Tasks and Datasets
We evaluate on continuous‑control D4RL MuJoCo tasks Fu2020D4RL . Throughout, we standardize states and actions using dataset statistics. When we interact with the environment we map actions back to the original scale. For each task, we standardize states/actions using dataset means/stds and adopt the D4RL raw return (higher is better).
We report three metrics. (i) Return: mean $\pm$ s.e. over $N_{\mathrm{roll}}$ episodes; (ii) Realized Type‑I: under $H_{0}$ with the deployed sampler, the frequency $\Pr[\ell_{\mathrm{cum}}\geq\hat{\tau}$] (target $\alpha$), shown in Figure 1; (iii) State‑conditional OOD: a kk‑NN proxy on $(s,a)$: for each generated pair, take the $k$ nearest in state and flag OOD if the action distance exceeds the qq‑th percentile of in‑dataset radii (defaults k=50k{=}50, $q{=}95\%$).
Our goal is to isolate the effect of inference-time guidance rather than to push absolute SOTA returns. Accordingly, we compare a standard $Q$-guided sampler (QG) against our LRT-guided sampler (LRT) and their simple composition (LRT+Q) under the same vanilla training pipeline, on representative D4RL MuJoCo tasks. This design controls for training confounders and highlights the structure of the sampler: does replacing a heuristic push by a calibrated, level-$\alpha$ gate improve the return–OOD trade-off? We therefore report raw return alongside a state-conditional OOD metric and realized Type-I, and interpret results through risk–performance curves and Pareto fronts rather than absolute leaderboards.
6.2 Baselines
  • LRT: evidence-gated sampler (no $Q$ step).
  • QG: standard action-space $Q$ update with schedules/clipping.
  • LRT+Q: LRT mean + small $Q$ step (calibration matches deployment).
6.3 Implementation & Reproducibility
Setup.
IQL critic ($\times2$56 MLP, $\gamma{=}0.99$, expectile 0.7); advantages on standardized $(s,a)$; labels: global top-$p$ (p=0.2p{=}0.2). Diffusion: T=50T{=}50, DDPM linear noise $(1{-}\alpha_{t})$ with endpoints $10^{-4}\!\to\!2{\times}10^{-2}$, MLP backbone (SiLU), two $\epsilon$ heads; AdamW $2{\times}10^{-4}$, batch 1024, 150 epochs. Class balancing via EMA positive rate; optional within-positive soft weights ($\tau_{A}$, $u_{\max}$). Full configs in Appx. E.
Inference & calibration.
Use reverse/posterior variance $\tilde{\sigma}_{t}^{2}$ for both heads and LLR; gate defaults $\beta_{\max}{=}1$, $\delta{\in}[1,2]$. Calibrate a single threshold $\hat{\tau}$ on a held-out state set by the fixed-point update
$\tau\leftarrow\mathrm{Quantile}_{1-\alpha}\{\ell_{\mathrm{cum}}(\tau)\}$
(K=6K{=}6 iterations with light momentum), then reuse $\hat{\tau}$ for all deployments with the same $(\beta_{\max},\delta)$ and $Q$-composition.
Protocol & compute.
Each configuration: $5$ seeds $\times$ 1010 episodes; report mean $\pm$ s.e.; actions mapped back to env scale and clipped; using D4RL-raw returns. Calibration uses $n\in[3000,5000]$ reverse trajectories.
6.4 Main Results
We sweep $\alpha\!\in\!\{0.20,0.10,0.05,0.02,0.01\}$ with fixed $(\beta_{\max},\delta)$ and compare QG, LRT, and LRT+Q. LRT is expected to (i) honor the target $\alpha$ (realized Type‑I $\approx\alpha$ within DKW tolerance), and (ii) trade return for lower state‑conditional OOD as $\alpha$ decreases, yielding improved Pareto fronts relative to QG in regimes where the critic is unreliable off‑support.
Table 1: Main results on D4RL MuJoCo. For each task/dataset and method, we report return (D4RL raw return) and state‑conditional OOD (%, lower is better). LRT honors the target $\alpha$ while improving the return–OOD trade‑off over QG in our implementation.
Data Name Mode Return (raw.) OOD
hopper-medium-replay-v2 lrt 377 0.017
hopper-medium-replay-v2 lrt+q 373 0.066
hopper-medium-replay-v2 q 366 0.023
halfcheetah-medium-replay-v2 lrt 578 0.011
halfcheetah-medium-replay-v2 lrt+q 786 0.094
halfcheetah-medium-replay-v2 q 706 0.092
walker2d-medium-replay-v2 lrt 362 0.003
walker2d-medium-replay-v2 lrt+q 436 0.033
walker2d-medium-replay-v2 q 433 0.032
Hopper
HalfCheetah
Walker2d
Figure 1: Risk–performance and Pareto fronts across tasks. Left of each row: risk–performance curves versus target $\alpha$. We report return (top), realized Type-I (middle), and state-conditional OOD (bottom) for LRT (solid) and QG (dashed); the realized Type-I tracks the target (gray) within finite-sample DKW bands. Right of each row: Pareto fronts (OOD vs. return; color encodes $\alpha$). LRT shifts the frontier up-and-left relative to QG on tasks where off-support critic error dominates, yielding higher return at lower OOD for the same $\alpha$. Error bars denote standard errors over evaluation rollouts.
Results at a glance.
Table 1 and Fig. 1 tell a consistent story. Across tasks, LRT achieves the lowest state-conditional OOD while keeping return competitive; adding a small value step (LRT+Q) raises return but at a higher OOD cost. Concretely, on Hopper LRT (377, 1.7%) outperforms QG (366, 2.3%) and avoids the OOD increase of LRT+Q (373, 6.6%); on HalfCheetah LRT+Q attains the top return (786, 9.4%) while LRT is conservative (578, 1.1%) and QG sits between (706, 9.2%); on Walker2d LRT minimizes OOD (0.3%), LRT+Q maximizes return (436, 3.3%), and QG is similar (433, 3.2%). Thus, LRT provides a low-risk anchor and the $Q$-composition moves the policy along the return–risk frontier.
Reading the curves and fronts.
In the left panels of Fig. 1, the realized Type-I (middle row) closely tracks the target $\alpha$ within DKW bands, validating calibration. As $\alpha$ decreases, return (top) drops modestly while state-conditional OOD (bottom) decreases consistently, exhibiting a smooth, monotone risk knob. The right panels (Pareto fronts) make the trade-off explicit: on Hopper and Walker2d, LRT shifts the frontier up-and-left relative to QG—higher return at lower OOD for a fixed $\alpha$—whereas on HalfCheetah, adding a small $Q$-step can push return further at the cost of OOD, matching the intended “anchor-plus-exploitation” behavior.
6.5 Ablations
Effect of $\alpha$.
Smaller $\alpha$ $\Rightarrow$ larger $\hat{\tau}$, fewer gate activations, lower state-conditional OOD with modest return drop (consistent with the left panels in Fig. 1).
Cap $\beta_{\max}$ & temperature $\delta$.
Decreasing $\beta_{\max}$ contracts updates toward $\mu_{u}$ (uniform OOD reduction even at large $\alpha$); decreasing $\delta$ sharpens the switch and approaches hard LRT (shifting Pareto up–left).
Labeling $p$.
Moderate $p\in[0.1,0.3]$ is robust; calibration preserves Type-I semantics regardless of $p$, so Pareto trends remain unchanged.
Evidence dynamics.
Prefix LLRs typically hover near zero at early high-noise steps and exceed $\hat{\tau}$ only later, if at all; the gate $\beta_{t}/\beta_{\max}$ remains near zero before crossing and then rises smoothly (Fig. 2). This matches the monotone risk knob seen in Fig. 1.
Figure 2: LLR traces. Prefix LLR across denoising steps on random states; lines are different trajectories. Data: hopper-medium-replay-v2.
7 Conclusion and Future Work
We presented LRT-Diffusion, a calibrated, inference-only guidance rule that turns diffusion-policy denoising into an evidence-gated process governed by a single, interpretable risk knob $\alpha$. Under equal covariances, the hard likelihood-ratio test (LRT) is uniformly most powerful at level $\alpha$; our smooth gate retains the same statistical semantics through a simple Monte-Carlo calibration matched to the deployed sampler. Empirically, LRT delivers calibrated control of Type-I risk and consistently improves the return–OOD trade-off over standard $Q$-guided sampling, while integrating seamlessly with a small critic step for additional exploitation. Together, these results demonstrate that risk-aware diffusion guidance can be achieved entirely at inference without modifying training objectives.
Several extensions naturally follow. (1) Learned or adaptive gating. Instead of a fixed calibrated threshold, the gate could be learned or amortized across tasks, or adapt a target $\alpha$ online under deployment constraints while preserving calibration. (2) Richer statistical models. Relaxing the equal-covariance assumption with heteroscedastic or learned head variances may yield stronger likelihood ratios and sharper decisions. (3) Sequential and early-stopping tests. Connecting the cumulative LLR to sequential or martingale tests (e.g., SPRT Wald1945SPRT ) would enable any-time risk control and early termination of denoising when the evidence is decisive, reducing compute without violating the level-$\alpha$ guarantee. (4) Broader integration and evaluation. Composing LRT with model-based planners or diffusion world models could link evidence-gated actions to trajectory-level reasoning, potentially higher returns on broader datasets.
Appendix A Gate Selection
A.1 Hard Gate-UMP: Proof of Proposition 1
Proof.
Define the (trajectory) likelihood ratio
$L(a_{0:T})\;:=\;\frac{p_{c}(a_{0:T}\mid s)}{p_{u}(a_{0:T}\mid s)}\;=\;\exp\!\big(\ell_{\mathrm{cum}}(a_{0:T})\big).$
Let $\Phi$ be the class of (possibly randomized) tests $\phi:\mathcal{A}^{T+1}\!\to[0,1]$ with size at most $\alpha$, i.e. $\mathbb{E}_{H_{0}}[\phi]\leq\alpha$. By the Neyman–Pearson lemma for simple $H_{0}$ vs. $H_{1}$, there exists $k>0$ such that the likelihood-ratio test $\phi^{\star}(a_{0:T})=\mathbf{1}\{L(a_{0:T})\geq k\}$ (with possible randomization on the boundary $\{L=k\}$) attains size exactly $\alpha$ and maximizes the power $\mathbb{E}_{H_{1}}[\phi]$ over $\Phi$. Since $\log(\cdot)$ is monotone, thresholding $L$ is equivalent to thresholding $\ell_{\mathrm{cum}}=\log L$. Hence the stated test is most powerful at level $\alpha$.
Remark. NP optimality holds for simple $H_{0}$ vs. $H_{1}$ regardless of covariance structure; the equal-covariance assumption is used only to compute $\ell_{t}$ in closed form.
A.2 Soft to Hard Limit: Proof of Lemma  1
Proof.
Recall from Sec. 4.3.2 the logistic gate
$\beta_{t}(\ell_{\mathrm{cum}};\tau,\delta,\beta_{\max})~=~\beta_{\max}\,\sigma\!\Big(\tfrac{\ell_{\mathrm{cum}}-\tau}{\delta}\Big),\qquad\sigma(x)=\tfrac{1}{1+e^{-x}},$
and the proposal mean at step $t$, $\mu_{t}^{\mathrm{soft}}=\mu_{u}+\beta_{t}(\mu_{c}-\mu_{u})$, with reverse variance $\sigma_{t}^{2}I$. We first establish the pointwise limit of $\beta_{t}$ and hence of $\mu_{t}^{\mathrm{soft}}$, then show weak convergence of the trajectory law by a coupling argument.
Step 1: Pointwise limit of the gate and the mean. Fix $(\tau,\beta_{\max})$. For any $x\in\mathbb{R}$ with $x\neq\tau$,
$\lim_{\delta\downarrow 0}\ \sigma\!\Big(\tfrac{x-\tau}{\delta}\Big)~=~\mathbf{1}\{x\geq\tau\}.$
Hence, for any realization of the running evidence $\ell_{\mathrm{cum}}$ at step $t$ such that $\ell_{\mathrm{cum}}\neq\tau$,
$\beta_{t}(\ell_{\mathrm{cum}};\tau,\delta,\beta_{\max})~\xrightarrow[\delta\downarrow 0]{}~\beta_{\max}\,\mathbf{1}\{\ell_{\mathrm{cum}}\geq\tau\}.$
Therefore
$\mu_{t}^{\mathrm{soft}}~=~\mu_{u}+\beta_{t}(\mu_{c}-\mu_{u})~\xrightarrow[\delta\downarrow 0]{}~\mu_{u}+\beta_{\max}\,\mathbf{1}\{\ell_{\mathrm{cum}}\geq\tau\}\,(\mu_{c}-\mu_{u})~=:\ \mu_{t}^{\mathrm{hard}}.$
It remains to note that, under our reverse kernels with Gaussian noise, the scalar $\ell_{\mathrm{cum}}$ is a continuous function of the Gaussian seeds, so $\mathbb{P}(\ell_{\mathrm{cum}}=\tau)=0$; hence the convergence above holds almost surely.
Step 2: Weak convergence of the trajectory law. Couple all samplers (for each $\delta\geq 0$) by using the same initial latent $a_{T}\sim q(\cdot)$ and the same Gaussian noise draws $z_{T},\ldots,z_{1}$:
$a_{t-1}^{\mathrm{soft}}\;=\;$ $\mu_{t}^{\mathrm{soft}}\big(t,s,a_{t}^{\mathrm{soft}}\big)\;+\;\sigma_{t}z_{t},$
$a_{t-1}^{\mathrm{hard}}\;=\;$ $\mu_{t}^{{\mathrm{hard}}}\big(t,s,a_{t}^{\mathrm{hard}}\big)\;+\;\sigma_{t}z_{t}.$
We prove by backward induction on $t=T,\ldots,1$ that $a_{t}^{\mathrm{soft}}\to a_{t}^{\mathrm{hard}}$ and the running evidence prefixes $\ell_{\mathrm{cum}}^{\mathrm{soft}}\to\ell_{\mathrm{cum}}^{\mathrm{hard}}$ almost surely as $\delta\downarrow 0$.
Base case $t=T$: By construction $a_{T}^{\mathrm{soft}}=a_{T}^{\mathrm{hard}}$ almost surely.
Inductive step: Assume $a_{t}^{\mathrm{soft}}\to a_{t}^{\mathrm{hard}}$ and $\ell_{\mathrm{cum}}^{\mathrm{soft}}\to\ell_{\mathrm{cum}}^{\mathrm{hard}}$ almost surely. The maps $(t,s,a)\mapsto\mu_{u}(t,s,a)$ and $\mu_{c}(t,s,a)$ are continuous (shared backbone + MLP heads with standard activations), hence
$\mu_{u}\big(t,s,a_{t}^{\mathrm{soft}}\big)\to\mu_{u}\big(t,s,a_{t}^{\mathrm{hard}}\big),\quad\mu_{c}\big(t,s,a_{t}^{\mathrm{soft}}\big)\to\mu_{c}\big(t,s,a_{t}^{\mathrm{hard}}\big)\quad\text{a.s.}$
By Step 1 and the a.s. convergence of $\ell_{\mathrm{cum}}^{\mathrm{soft}}$,
$\beta_{t}^{\mathrm{soft}}=\beta_{t}(\ell_{\mathrm{cum}}^{\mathrm{soft}};\tau,\delta,\beta_{\max})\to\beta_{\max}\mathbf{1}\{\ell_{\mathrm{cum}}^{\mathrm{hard}}\geq\tau\}$
almost surely (the boundary has probability zero). Therefore
$\mu_{t}^{\mathrm{soft},(\delta)}~\longrightarrow~\mu_{t}^{\mathrm{hard}}\quad\text{a.s.}$
With the coupled noise $z_{t}$, we obtain
$a_{t-1}^{\mathrm{soft}}=\mu_{t}^{\mathrm{soft}}+\sigma_{t}z_{t}~\longrightarrow~\mu_{t}^{\mathrm{hard}}+\sigma_{t}z_{t}=a_{t-1}^{\mathrm{hard}}\quad\text{a.s.}$
Finally, the one-step LLR increment $\ell_{t}$ is a continuous function of $(a_{t},a_{t-1})$ under equal covariances, hence $\ell_{t}^{\mathrm{soft}}\to\ell_{t}^{\mathrm{hard}}$ and the updated evidence $\ell_{\mathrm{cum}}^{\mathrm{soft}}\to\ell_{\mathrm{cum}}^{\mathrm{hard}}$ almost surely, closing the induction.
Thus $a_{0:T}^{\mathrm{soft}}\to a_{0:T}^{\mathrm{hard}}$ almost surely under the coupling. Almost-sure convergence under a coupling implies weak convergence of the induced trajectory laws, so the sampler with soft gate converges (in distribution) to the sampler with the hard gate as $\delta\downarrow 0$. ∎
Appendix B Proofs for Finite-sample Calibration and Stability
Frozen sampler and notation.
Fix the deployed (frozen) sampler, including $(\beta_{\max},\delta)$ and, if enabled, the deterministic $Q$-composition (schedule, clipping, and the evaluation center). Under $H_{0}$ the trajectory $a_{T:0}$ is generated by this sampler using the unconditional head. Let $F_{0}$ be the CDF of the cumulative LLR $\ell_{\mathrm{cum}}$ under $H_{0}$, and let $\widehat{F}_{n}$ be the empirical CDF from $n$ i.i.d. calibration runs. The plug-in threshold is $\hat{\tau}=\inf\{x:\widehat{F}_{n}(x)\geq 1-\alpha\}$.
B.1 Finite-sample Calibration: Proof of Theorem 2
Proof.
Let $F_{0}$ be the CDF of the cumulative LLR $\ell_{\mathrm{cum}}$ under $H_{0}$ for the frozen sampler (soft gate and, if present, the fixed $Q$-step). Given i.i.d. calibration draws $\ell^{(1)},\dots,\ell^{(n)}$ and the empirical CDF $\widehat{F}_{n}$, define the plug-in quantile $\hat{\tau}=\inf\{x:\widehat{F}_{n}(x)\geq 1-\alpha\}$.
According to Dvoretzky–Kiefer–Wolfowitz
$\mathbb{P}r\left(\sup_{x}|F_{0}(x)-\widehat{F}_{n}(x)|\geq\varepsilon_{n}\right)\leq 2e^{-2n\epsilon_{n}^{2}}$ (13)
Let $\zeta=2e^{-2n\epsilon_{n}^{2}}$ denote the lower bound for eq (13). Then,
$\mathbb{P}r\left(\sup_{x}|F_{0}(x)-\widehat{F}_{n}(x)|\leq\varepsilon_{n}\right)\geq 1-2e^{-2n\epsilon_{n}^{2}}=1-\zeta$ (14)
Hence, $F_{0}(\hat{\tau})\geq\widehat{F}_{n}(\hat{\tau})-\varepsilon_{n}\geq 1-\alpha-\varepsilon_{n}$ and $\mathbb{P}_{H_{0}}(\ell_{\mathrm{cum}}\geq\hat{\tau})\leq\alpha+\varepsilon_{n}$. Moreover, when $1-2e^{-2n\epsilon_{n}^{2}}\geq 1-\zeta$ ensure at least $1-\zeta$ that the prediction is accurate, which the sample size will follow $n\;\geq\;\frac{1}{2\varepsilon^{2}}\,\log\!\frac{2}{\zeta}$. ∎
B.2 Deterministic Displacement bound: Proof for Lemma 3
Proof.
By definition of the (deterministic) update at step $t$,
$a_{t}-\mu_{u,t}=\beta_{t}\,\Delta\mu_{t}\;+\;\lambda_{t}\,\sigma_{t}^{2}\,g_{t}.$
Taking Euclidean norms and applying the triangle inequality,
$\|a_{t}-\mu_{u,t}\|\;\leq\;|\beta_{t}|\,\|\Delta\mu_{t}\|\;+\;|\lambda_{t}|\,\sigma_{t}^{2}\,\|g_{t}\|.$
Using the bounds $0\leq\beta_{t}\leq\beta_{\max}$, $0\leq\lambda_{t}\leq\lambda_{\max}$ and $\|g_{t}\|\leq G$, we obtain
$\|a_{t}-\mu_{u,t}\|\;\leq\;\beta_{\max}\,\|\Delta\mu_{t}\|\;+\;\lambda_{\max}\,\sigma_{t}^{2}\,G\;=:\;B_{t},$
which is (12).
If, in addition, $\|\Delta\mu_{t}\|\leq D$ and $\sigma_{t}^{2}\leq S^{2}$, then
$\|a_{t}-\mu_{u,t}\|\;\leq\;\beta_{\max}D\;+\;\lambda_{\max}S^{2}G\;=:\;B_{\text{step}}\quad\text{for all }t.$
Finally, summing the per-step bounds yields the stated cumulative bound:
$\sum_{t=1}^{T}\|a_{t}-\mu_{u,t}\|\;\leq\;\sum_{t=1}^{T}B_{t}\;\leq\;T\,B_{\text{step}}.$
B.3 One-step LLR Decomposition and Moment Bounds
Under equal covariances $\sigma_{t}^{2}I$, the one-step LLR admits the exact identity
$\Delta\ell_{t}=-\frac{\|a_{t-1}-\mu_{c,t}\|^{2}-\|a_{t-1}-\mu_{u,t}\|^{2}}{2\sigma_{t}^{2}}=\frac{\Delta\mu_{t}^{\top}\!\left(a_{t-1}-\tfrac{\mu_{c,t}+\mu_{u,t}}{2}\right)}{\sigma_{t}^{2}}.$ (15)
Write the stochastic update as $a_{t-1}=m_{t}+\sigma_{t}z_{t}$, where $z_{t}\sim\mathcal{N}(0,I)$ is independent of $z_{t^{\prime}}$ for $t^{\prime}\neq t$ and independent of the past given the sampler’s history (the filtration $\mathcal{F}_{t}$ generated by $(a_{T},\ldots,a_{t})$ and the gating/gradient choices). Substituting into (15) yields the deterministic + Gaussian decomposition
$\Delta\ell_{t}=\underbrace{\frac{\Delta\mu_{t}^{\top}\!\left(m_{t}-\tfrac{\mu_{c,t}+\mu_{u,t}}{2}\right)}{\sigma_{t}^{2}}}_{=:~A_{t}\text{ (deterministic given }\mathcal{F}_{t})}\;+\;\underbrace{\frac{\Delta\mu_{t}^{\top}z_{t}}{\sigma_{t}}}_{=:~B_{t}\text{ (zero-mean Gaussian)}}.$
(i) Conditional mean bound.) Since $m_{t}-\tfrac{\mu_{c,t}+\mu_{u,t}}{2}=(m_{t}-\mu_{u,t})-\tfrac{1}{2}\Delta\mu_{t}$, by Cauchy–Schwarz,
$|A_{t}|\;=\;\frac{1}{\sigma_{t}^{2}}\,\big|\langle\Delta\mu_{t},\,(m_{t}-\mu_{u,t})-\tfrac{1}{2}\Delta\mu_{t}\rangle\big|\ \leq\ \frac{\|\Delta\mu_{t}\|}{\sigma_{t}^{2}}\Big(\|m_{t}-\mu_{u,t}\|+\tfrac{1}{2}\|\Delta\mu_{t}\|\Big).$
Taking expectation conditional on $\mathcal{F}_{t}$ gives
$\big|\mathbb{E}[\Delta\ell_{t}\mid\mathcal{F}_{t}]\big|\ \leq\ \frac{\|\Delta\mu_{t}\|}{\sigma_{t}^{2}}\Big(\|m_{t}-\mu_{u,t}\|+\tfrac{1}{2}\|\Delta\mu_{t}\|\Big)\ \leq\ \frac{\|\Delta\mu_{t}\|}{\sigma_{t}^{2}}\Big(B_{t}+\tfrac{1}{2}\|\Delta\mu_{t}\|\Big),$
where in the last inequality we used the displacement bound $\|m_{t}-\mu_{u,t}\|\leq B_{t}$ from §B.2.
(ii) Conditional variance.) Because $B_{t}=\frac{1}{\sigma_{t}}\Delta\mu_{t}^{\top}z_{t}$ with $z_{t}\sim\mathcal{N}(0,I)$,
$\mathrm{Var}(\Delta\ell_{t}\mid\mathcal{F}_{t})=\mathrm{Var}(B_{t}\mid\mathcal{F}_{t})=\frac{\|\Delta\mu_{t}\|^{2}}{\sigma_{t}^{2}}.$
In particular, if $\|\Delta\mu_{t}\|\leq D$ and $\sigma_{t}\geq\sigma_{\min}>0$, then $\mathrm{Var}(\Delta\ell_{t}\mid\mathcal{F}_{t})\leq D^{2}/\sigma_{\min}^{2}$ and
$\big|\mathbb{E}[\Delta\ell_{t}\mid\mathcal{F}_{t}]\big|\ \leq\ \frac{D}{\sigma_{\min}^{2}}\Big(B_{\text{step}}+\tfrac{D}{2}\Big)\qquad\text{(using }\|m_{t}-\mu_{u,t}\|\leq B_{\text{step}}\text{).}$
B.4 Sub-Gaussian concentration for the cumulative LLR
Define the variance proxy
$V\;:=\;\sum_{t=1}^{T}\frac{\|\Delta\mu_{t}\|^{2}}{\sigma_{t}^{2}}.$
Let $\mathcal{F}_{t}$ be the sampler’s filtration up to step $t$. As above, given $\mathcal{F}_{t}$, we can write $\Delta\ell_{t}=A_{t}+B_{t}$ with $B_{t}=\frac{1}{\sigma_{t}}\Delta\mu_{t}^{\top}z_{t}$, where $z_{t}\sim\mathcal{N}(0,I)$ is independent of $\mathcal{F}_{t}$ and of $z_{t^{\prime}}$ for $t^{\prime}\neq t$. Hence, conditionally,
$\mathbb{E}\big[e^{\lambda B_{t}}\mid\mathcal{F}_{t}\big]=\exp\!\left(\frac{\lambda^{2}}{2}\cdot\frac{\|\Delta\mu_{t}\|^{2}}{\sigma_{t}^{2}}\right).$
Iterating conditional expectations (tower property) and using independence of the $z_{t}$ across $t$,
$\mathbb{E}\!\left[\exp\!\left(\lambda\sum_{t=1}^{T}(B_{t}-\mathbb{E}[B_{t}\mid\mathcal{F}_{t}])\right)\right]\leq\exp\!\left(\frac{\lambda^{2}}{2}\sum_{t=1}^{T}\frac{\|\Delta\mu_{t}\|^{2}}{\sigma_{t}^{2}}\right)=\exp\!\left(\frac{\lambda^{2}}{2}V\right).$
Therefore $\sum_{t=1}^{T}\big(B_{t}-\mathbb{E}[B_{t}\mid\mathcal{F}_{t}]\big)$ is sub-Gaussian with proxy $V$. Adding the deterministic (given history) terms $A_{t}-\mathbb{E}[A_{t}\mid\mathcal{F}_{t}]$ only shifts the mean and does not affect the sub-Gaussian proxy. Consequently,
$\ell_{\mathrm{cum}}-\mathbb{E}[\ell_{\mathrm{cum}}]=\sum_{t=1}^{T}\big(\Delta\ell_{t}-\mathbb{E}[\Delta\ell_{t}\mid\mathcal{F}_{t}]\big),$
is sub-Gaussian with variance proxy V, and for any $x>0$,
$\Pr\!\left[\ell_{\mathrm{cum}}-\mathbb{E}\ell_{\mathrm{cum}}\geq x\right]\ \leq\ \exp\!\left(-\frac{x^{2}}{2V}\right).$
Under the scale controls in §B.2,
$V\leq\sum_{t=1}^{T}D^{2}/\sigma_{\min}^{2}=T\,D^{2}/\sigma_{\min}^{2}$
Summary of implications.
The cap $\beta_{\max}$ and gradient clipping (via $G,\lambda_{\max}$) yield a per-step deterministic displacement $B_{t}$ and a uniform bound $B_{\text{step}}$; together with the equal-covariance identity, these give (i) explicit bounds on the conditional mean and variance of each LLR increment, and (ii) a sub-Gaussian concentration for the cumulative evidence with proxy $V=\sum_{t}\|\Delta\mu_{t}\|^{2}/\sigma_{t}^{2}$. These stability controls make the calibration reproducible (Theorem 2) and the gate numerically well-conditioned.
Appendix C When LRT Dominate Q guidance
C.1 Proof for Proposition  3
Proof.
Write $\Lambda(s,a):=\hat{Q}(s,a)-Q^{\rm*}(s,a)$. Then, under the assumption in Sec 5.3, under policy $\pi$ we have
$|\Lambda(s,a_{\pi})|\leq\epsilon_{in}(1-\eta_{\pi})+\epsilon_{out}\eta_{\pi}.$
Lower bound for $\mathbb{E}_{s}[Q^{\rm*}(s,a_{\rm LRT})]$:
$\mathbb{E}_{s}[Q^{\rm*}(s,a_{\rm LRT})]=$ $\mathbb{E}_{s}[\hat{Q}(s,a_{LRT})-\hat{Q}(s,a_{LRT})+Q^{\rm*}(s,a_{\rm LRT})]$ (16)
$\geq$ $E_{s}[\hat{Q}(s,a_{LRT})]-\Lambda(s,a_{LRT})$ (17)
$\geq$ $E_{s}[\hat{Q}(s,a_{LRT})]-|\Lambda(s,a_{LRT})|$ (18)
$\geq$ $E_{s}[\hat{Q}(s,a_{LRT})]-\epsilon_{in}-(\epsilon_{out}-\epsilon_{in})\eta_{LRT}.$ (19)
Similarly, the upper bound for $\mathbb{E}_{s}[Q^{\rm*}(s,a_{\rm Q})]$:
$\mathbb{E}_{s}[Q^{\rm*}(s,a_{\rm Q})]=$ $\mathbb{E}_{s}[\hat{Q}(s,a_{Q})-\hat{Q}(s,a_{Q})+Q^{\rm*}(s,a_{\rm Q})]$ (20)
$\leq$ $E_{s}[\hat{Q}(s,a_{Q})]+|\Lambda(s,a_{Q})|$ (21)
$\leq$ $E_{s}[\hat{Q}(s,a_{Q})]+\epsilon_{in}+(\epsilon_{out}-\epsilon_{in})\eta_{Q}.$ (22)
Therefore
$\mathbb{E}_{s}[Q^{\rm true}(s,a_{\rm LRT})]\ \geq\ \mathbb{E}_{s}[\hat{Q}(s,a_{\rm LRT})]-\varepsilon_{\rm in}-\nu\,\eta(\pi_{\rm LRT}),$
$\mathbb{E}_{s}[Q^{\rm true}(s,a_{Q})]\ \leq\ \mathbb{E}_{s}[\hat{Q}(s,a_{Q})]+\varepsilon_{\rm in}+\nu\,\eta(\pi_{Q}).$
Subtract the two displays and rearrange to obtain the stated bound for $\Delta_{Q^{\rm true}}$.
$\Delta_{Q^{true}}\geq\Delta_{\hat{Q}}-2\varepsilon_{\rm in}-\nu\,\left(\eta(\pi_{Q})+\eta(\pi_{\rm LRT})\right)$ (23)
C.2 Condition for LRT to Dominate
Corollary 3.1 (Feasible $\alpha$ via the union bound).
Under Assumption 4 with $c(\alpha)\leq\alpha T$, a sufficient condition for $\Delta_{\text{true}}\geq 0$ is
$\alpha~\leq~\alpha_{\max}:=\frac{\Delta_{\widehat{Q}}-2\varepsilon_{\rm in}-\Delta_{\varepsilon}\,\eta(\pi_{Q})}{\Delta_{\varepsilon}\,T}.$ (24)
In particular, the feasible set $\{\alpha:\Delta_{\text{true}}\geq 0\}$ is nonempty whenever $\Delta_{\widehat{Q}}>2\varepsilon_{\rm in}+\Delta_{\varepsilon}\,\eta(\pi_{Q})$.
Proof.
Plug $\eta(\pi_{\rm LRT})\leq\alpha T$ into (23) and solve $\Delta_{\text{true}}\geq 0$ for $\alpha$. Since $\Delta_{\varepsilon}T>0$ unless $T=0$ or $\Delta_{\varepsilon}=0$, (24) follows. ∎
How to widen the feasible interval.
From (24), the numerator increases if we (i) enlarge $\Delta_{\widehat{Q}}$ (better critic or stronger evidence-tied guidance), and/or (ii) reduce $\eta(\pi_{Q})$ (conservative anchoring/clipping for Q-steps). Moreover, taking smaller $\alpha$ or smaller $\beta_{\max}$ further reduces $\eta(\pi_{\rm LRT})$.
C.3 A tighter, Data-driven Alternative
The surrogate $\eta(\pi_{\rm LRT})\leq\alpha T$ is often loose. Calibrate once and measure $\widehat{\eta}_{\rm LRT}(\alpha)$ on held-out states (our kNN proxy). Then the same derivation yields the data-driven bound
$\Delta_{\text{true}}~\geq~\Delta_{\widehat{Q}}~-~2\varepsilon_{\rm in}~-~\Delta_{\varepsilon}\big(\widehat{\eta}_{\rm LRT}(\alpha)+\widehat{\eta}(\pi_{Q})\big),$ (25)
so any $\alpha$ satisfying $\Delta_{\widehat{Q}}\geq 2\varepsilon_{\rm in}+\Delta_{\varepsilon}\big(\widehat{\eta}_{\rm LRT}(\alpha)+\widehat{\eta}(\pi_{Q})\big)$ is sufficient. DKW-type concentration gives a uniform band for the empirical OOD curve, translating into a high-probability guarantee for (25).
C.4 Optional Refinement with Consistency or Distance-based Bounds
If the critic is consistent on support, e.g., with probability $1-\delta_{n}$, $\varepsilon_{\rm in}\leq r_{n}\!\to\!0$ as $n\to\infty$, then
$\Delta_{\text{true}}~\geq~\Delta_{\widehat{Q}}~-~2r_{n}~-~\Delta_{\varepsilon}\big(\eta(\pi_{\rm LRT})+\eta(\pi_{Q})\big),$
loosening the requirement on $\Delta_{\widehat{Q}}$ as $n$ grows. Alternatively, if there exists $L_{Q}$ such that $|\widehat{Q}-Q^{\text{true}}|\leq\varepsilon_{\rm in}+L_{Q}\,\mathrm{dist}(a,\mathcal{S}(s))$, then
$\Delta_{\text{true}}~\geq~\Delta_{\widehat{Q}}-2\varepsilon_{\rm in}-L_{Q}\,\mathbb{E}_{s}\!\big[\mathrm{dist}(a_{\rm LRT},\mathcal{S}(s))+\mathrm{dist}(a_{Q},\mathcal{S}(s))\big].$
Using our per-step movement bound and the monotonic effect of $\alpha,\beta_{\max}$ on the state-conditional distance, the last expectation can be turned into an explicit function of $(\alpha,\beta_{\max})$, yielding a more problem-adapted feasible range for $\alpha$.
C.5 Further Discussion for LRT dominate Q
Combining Prop. 3 and Prop. 4 yields the sufficient condition
$\Delta_{Q^{\rm true}}~\geq~\Delta_{\hat{Q}}~-~2\varepsilon_{\rm in}~-~\kappa\big(\alpha T+\eta(\pi_{Q})\big),$
$\alpha_{\max}~=~\frac{\Delta_{\hat{Q}}-2\varepsilon_{\rm in}-\kappa\,\eta(\pi_{Q})}{\kappa\,T},$
where
$\Delta_{Q^{\rm true}}:=\mathbb{E}_{s}\!\left[Q^{\rm true}(s,a_{\rm LRT})-Q^{\rm true}(s,a_{Q})\right],$
and
$\Delta_{\hat{Q}}:=\mathbb{E}_{s}\!\left[\hat{Q}(s,a_{\rm LRT})-\hat{Q}(s,a_{Q})\right].$
The bound decomposes into three interpretable pieces:
  • Predicted–value gap $\Delta_{\hat{Q}}$. How much the critic $\hat{Q}$ prefers LRT samples over Q–guided samples on average. If the critic systematically overestimates off–support actions, $\Delta_{\hat{Q}}$ can be small or negative; the condition is sufficient only and may fail even when LRT is better in the environment.
  • On–support error $2\varepsilon_{\rm in}$. A safety margin independent of $\alpha$. Better–fit (or pessimistic/trimmed) critics reduce this term and enlarge the feasible range.
  • OOD penalty $\kappa(\alpha T+\eta(\pi_{Q}))$. The LRT part scales with $\alpha$ and with the number of gated steps $T$ via the union proxy $\eta(\pi_{\rm LRT})\leq\alpha T$; the Q–guided baseline pays $\kappa\,\eta(\pi_{Q})$. Smaller $\kappa$—e.g., via a smaller $\beta_{\max}$ or late–step gating when $\sigma_{t}$ is small—makes the condition easier to satisfy.
C.5.1 When does LRT “dominate” Q–guidance?
If $\alpha\leq\alpha_{\max}$, then $\mathbb{E}_{s}[Q^{\rm true}(s,a_{\rm LRT})]\geq\mathbb{E}_{s}[Q^{\rm true}(s,a_{Q})]$ is certified. Two practical corollaries are useful:
  • A small–OOD Q baseline helps the certificate. If $\eta(\pi_{Q})$ is small (Q–guidance rarely leaves support), the denominator in $\alpha_{\max}$ is fixed while the numerator increases, so $\alpha_{\max}$ grows: $\eta(\pi_{Q})\!\downarrow\Rightarrow\alpha_{\max}\!\uparrow$. In the extreme $\eta(\pi_{Q}){=}0$, the feasible range reads $\alpha\leq(\Delta_{\hat{Q}}-2\varepsilon_{\rm in})/(\kappa T)$. Intuitively, when Q–guidance already behaves conservatively, we can open the LRT gate more (choose a larger $\alpha$) and still retain a certificate.
  • Levers that enlarge $\alpha_{\max}$. (i) Reduce $T$ by gating only in late steps (then $\alpha_{\max}\propto 1/T$); (ii) Lower $\kappa$ (e.g., smaller $\beta_{\max}$ or restricting gating to small–variance steps; anchoring gradients at $\mu_{u}$ also lowers off–support effects); (iii) Increase $\Delta_{\hat{Q}}$ by taking a modest Q–step near $\mu_{\rm LRT}$ when the gate opens or by using evidence–tied/annealed guidance.
C.5.2 Sweeping $\alpha$ in practice.
The certificate is conservative for two reasons: (i) it uses the union proxy $\eta(\pi_{\rm LRT})\!\leq\!\alpha T$; (ii) $\Delta_{\hat{Q}}$ can be biased when $\hat{Q}$ overestimates OOD. Thus even when $\alpha>\alpha_{\max}$ and the inequality cannot be certified, LRT may still outperform Q–guidance empirically. In experiments we therefore sweep $\alpha$ on a logarithmic grid and select the knee of the return–risk curve (Fig.  1); the bound serves as a sanity upper bound for $\alpha$, not as a hard constraint.
A practical process.
  1. 1.
    On a small held–out state batch, estimate $\widehat{\Delta}_{\hat{Q}}$ and $\widehat{\eta}(\pi_{Q})$ by generating actions from both samplers and evaluating $\hat{Q}$ and the state–conditional OOD rate.
  2. 2.
    Form the diagnostic
    $\widehat{\alpha}_{\max}~=~\frac{\widehat{\Delta}_{\hat{Q}}-2\varepsilon_{\rm in}-\kappa\,\widehat{\eta}(\pi_{Q})}{\kappa\,T}\quad$
    (use a conservative constant if $\varepsilon_{\rm in}$ is unknown).
  3. 3.
    Sweep $\alpha$ on a log grid in $\min(0.5\,\widehat{\alpha}_{\max},\,0.2),\,0.2\,]$, and select the knee on the return–risk curves using the calibrated LRT sampler.
Why $T$ matters.
For T=50T{=}50 and $\kappa{\approx}1$, even $\Delta_{\hat{Q}}-2\varepsilon_{\rm in}\approx 0.4$ gives $\alpha_{\max}\approx 0.008$ (quite small). Gating only in the last $10$ steps effectively sets T=10T{=}10, yielding $\alpha_{\max}\approx 0.04$. This explains why late–step gating / smaller $\beta_{\max}$ are both stable and easier to certify in practice.
Appendix D Appending with Q-guidance
D.1 Choice of the evaluation center $a_{c}$.
  • $a_{c}=\mu_{u}$ (conservative). Evaluate $\nabla_{a}\hat{Q}$ near the behavior-supported mean; most stable and robust off-support.
  • $a_{c}=\mu_{\text{LRT},t}$ (exploitative). When evidence opens the gate ($\beta_{t}$ large), probe gradients closer to the “good” direction.
  • $a_{c}=(1-\rho)\mu_{u}+\rho\,\mu_{\text{LRT},t}$ (interpolate). $\rho\!\in[0,1]$ trades conservatism for exploitation; a simple adaptive choice is $\rho=\beta_{t}/\beta_{\max}$.
Appendix E Implementation Details
Training. IQL critic: MLPs (two hidden layers, 256 units), $\gamma\!=\!0.99$, expectile $0.7$ (following IQL [7]), Adam ($3{\times}10^{-5}$), batch 1024, $30$ epochs. Advantages are computed on standardized $(s,a)$. Labels: global top‑$p$ with p=0.2p\!=\!0.2 unless stated. Diffusion: T=50T\!=\!50, DDPM linear schedule on the forward noise $(1-\alpha_{t})$ with endpoints $(1-\alpha_{1}\!=\!10^{-4},\,1-\alpha_{T}\!=\!2{\times}10^{-2})$, , backbone MLP with SiLU activations, two $\epsilon$‑heads; AdamW ($2{\times}10^{-4}$), batch 1024, 150 epochs. Class‑balancing with an EMA of the positive rate; optional soft advantage weight within positives with temperature $\tau_{A}$ and cap $u_{\max}$.
Inference and calibration. We use the posterior variance $\tilde{\sigma}_{t}^{2}$ (DDPM reverse variance) for both heads and LLR; alternative variance parameterizations have also been explored [21]. Gate defaults: $\beta_{\max}\!=\!1$, $\delta\!\in\![1,2]$ (ablate). We calibrate $\hat{\tau}$ on a held‑out state set (disjoint from any rollout episodes) with a fixed‑point update
$\tau\leftarrow\mathrm{Quantile}_{1-\alpha}\big\{\ell_{\mathrm{cum}}(\tau)\big\},$
running $K$ iterations (default K=6K\!=\!6) with light momentum to avoid oscillation, and then reuse the resulting $\hat{\tau}$ for all deployments with the same $(\beta_{\max},\delta)$ and (if enabled) the same QQ‑composition.
Evaluation protocol. Unless stated, each configuration is evaluated over $5$ random seeds $\times$ 1010 episodes per seed. We report means $\pm$ standard errors across all episodes. Actions are mapped back to the environment scale and clipped to bounds before stepping. For D4RL environments we also report normalized scores (when supported).